NashaBeMe:

What is 5 times 5 to the third power

NashaBeMe:

thank u guys

iHeLpYoUuHeLpMe:

Ik this by his tells you to multiply 1 by the number as many times as the exponent says. The example above is 2 raised to the third power (raised to the third power means the exponent is 3). This is equivalent to the multiplication problem below because there is a 1 multiplied by 2 three times.

NashaBeMe:

i wanted to give yall both best response how do i

iHeLpYoUuHeLpMe:

You can just have one person

jhonyy9:

(5*5)^3 = ?

jhonyy9:

or 5*5^3 = ?

KiWiPoP1:

Yeah youre gonna need to clarify, which one of those ^ is it

ramen:

\[(5*5)^3\] = 5*5 which is 25. and what is \[25^3\] you can input it into a calculator.

koolkidjay:

\[25^{3}= 625\]

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @koolkidjay \[25^{3}= 625\] \(\color{#0cbb34}{\text{End of Quote}}\) no 625 = 25^2